A rectangular blanket is 20 inches longer than it is wide. If the area of the blanket is 1,125 square inches, find the length and the width.
Answer by tutorcecilia(2152)

A rectangular blanket is 20 inches longer than it is wide. If the area of the blanket is 1,125 square inches, find the length and the width.
A=(length)(width)
A=1125
Length=w+20
Width=w
.
1125=(w+20)(w) [plug-in the values]
1125=w^2+20w
0=w^2+20w-1125
0=(w-25)(w+45)[factor and solve for the w-term]
w-25=0
w=25 inches [this is the solution for the width]
and
w+45=0
w=-45 [eliminate this solution since a measurement cannot be negative]
.
Since the length=w+20, than, 25+20=45 inches (length)